• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 08:07:25 on February 9,2002

    Entry number 82350

    keyword=Owl Shift summary

    Begin shift with 5D kinematics.

    23:56 - 00:26 Run # 2314 LH2, Norad. mon, veto off

    00:30 Move to LH2 + rad. target, move spectrometer angle to 28.91 degrees.
    00:35 Change Left spectrometer momentum to 3.154 GeV/c
    00:42 Asked for beam to check prescale factors.

    00:48 - 01:08 Run # 2315 LH2 + rad. mon, veto off
    01:09 - 01:26 Run # 2316 LH2 + rad, mon, veto off
    01:27 - 01:43 Run # 2317 LH2 + rad, mon, veto off

    Coda problems runs 2318 and 2319 junk

    01:51 - 02:08 Run # 2320 LH2 + rad, mon, veto off
    02:08 - 02:20 Run # 2321 LH2 + rad, mon, veto off
    During run 2321 Coda crashed.

    02:15 Gas alarm went off for Argon. Level at -3.5 PSI
    02:22 Dipangkar went to check the gas shed
    02:30 Talked to Xiaodong, and he said to change the Argon bottle and continue taking data with the current situation.

    02:31 - 2:50 Run # 2322 LH2 + rad, mon, veto off, low Argon pressure.

    02:50 VDC's have not recovered. Paged Jack Segal, and he is looking into the problem.

    03:08 Turned off HV for FPP.
    03:20 Jack came by, and he reset the switch for the main relay on the gas panel in the gas shed. The FPP gas flow has come back up, and the VDC gas flow is slowly coming up too.

    03:35 More Coda problems. kcoda and restart. Run # 2323 is junk
    03:38 - 03:55 Run # 2324 LH2 + rad, mon, veto off
    03:55 - 04:14 Run # 2325 LH2 + rad, mon, veto off, end of run failed: TS2 reboot

    04:05 Calorimeter HV trip

    04:20 - 04:41 Run # 2326 LH2 + rad, mon, veto off
    04:41 - 04:59 Run # 2327 LH2 + rad, mon, veto off
    05:00 - 05:16 Run # 2328 LH2 + rad, mon, veto off
    05:17 - 05:38 Run # 2329 LH2 + rad, mon, veto off
    05:38 - 05:55 Run # 2330 LH2 + rad, mon, veto off
    05:56 - 06:16 Run # 2331 LH2 + rad, mon, veto off
    06:16 - 06:36 Run # 2332 LH2 + rad, mon, veto off
    06:37 - 06:54 Run # 2333 LH2 + rad, mon, veto off, end of run failed

    06:58 Called MCC to stop delivering beam to turn the veto on. One HV channel on the veto (# 16) did not want to set. Igor reset it manually. (Halog entry: 82343)

    07:05 Bogdan called and wanted us to run with veto and magnet off to check the data, before running with veto and magnet on.
    07:08 Magnet and veto off.

    07:12 - 07:26 Run # 2334 LH2 + rad, moff, veto off. Stopped for QE measurement.
    07:33 - 07:53 Run # 2335 LH2 + rad, moff, veto off
    Total: ~ 0.04 C for 5D(f)

    07:55 Call MCC to take beam away to turn magnet and veto on
    07:57 Asked for 5 uA of rastered beam. Current at 4 uA.
    08:05 Start run # 2336 LH2 + rad, mon, veto on